OpenStudy (anonymous):

solve the following log equation ln (x+2) - ln (x-4) = 5

OpenStudy (blockcolder):

ln((x+2)/(x-4))=5 e^5=(x+2)/(x-4) (x-4)e^5=x+2 (e^5)x-4e^5=x+2 (e^5)x-x=2+4e^5 x(e^5-1)=2+4e^5 x=(2+4e^5)/(e^5-1)